# Distributional Scaling of Emergent Capabilities

### v3

@@ 4.1 Emergence Across Data Compositions
We first examine how data composition influences the emergence of bimodal behavior in Qwen2.5-0.5B. Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s bimodal performance distribution in ways that mirror the effect of scale.
In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 [34], most models (out of 20 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format [13]. As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline.
@@ 4.2 Emergence Across Model Scales
Next, we show that, as in the synthetic setting, LM model size affects the formation of performance clusters. Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 7 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models [35, 21]. In Appendix E.1, we show similar results in the Commonsense QA [43] task.
5 Related Works
Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order. Prior studies have documented performance differences across various stress test sets [4, 30]. Such differences persist throughout training, not just at the final checkpoint [50]. Dodge et al. [6] compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance. Additionally, Sellam et al. [38] demonstrated that LMs independently pre-trained from scratch can differ in downstream task performance and social biases. Existing work has also found model runs can cluster in out-of-distribution behavior [19] and in training dynamics [33, 14], hinting at multimodal variation. Our work connects these random clustering effects to the phenomenon of emergence at scale.
Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on [1, 5, 9, 15, 37, 48]. Various proposals have, consequently, aimed to improve length generalization. Some methods focus on alternative positional encodings [39, 32, 42, 22, 17]. Others modify the dataset format by adding scratchpad or Chain-of-Thought [1] or by incorporating padding and index hints [16, 49]. Regarding random variation, Zhou et al. [49] and Zhou et al. [51] provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
Emergent abilities of LMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets [12, 35, 2, 21]. These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales [45, 8, 41]. Understanding the conditions and mechanisms underlying emergence is a key area of research. Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale [41, 36], some breakthrough capabilities remain emergent. Snell et al. [40] found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size [44]. For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model [31]. Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. [7] show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. [24] use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
6 Discussion and Conclusions
@@
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale. Previous work has already documented the variability of length generalization across random seeds [51, 49]; in general, out-of-distribution behavior like compositional rules [27] or associative biases [38] often exhibit extreme variation compared to in-distribution performance. We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics [19] and even in the timing of generalization breakthroughs during training [14]. The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 3.3); the transition from the initial unimodal distribution to a bimodal one is sudden.
In the wider scaling laws literature, which often focuses on aggregated loss, parameters added by increasing depth and width are often treated interchangeably in their contribution to total compute. By separately considering depth and width, our results suggest that emergent phenomena may be sensitive to tradeoffs between architectural hyperparameters. In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs.
@@ Appendix A Experimental details on synthetic tasks
Below we provide more details when training decoder-only language models on the count and addition—with index hints— tasks. Hyperparameters are largely adapted from [49]. We train all of our models to convergence on the train distribution. Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
@@
  • Model scales: As mentioned in Section 2, we scale up our models by fixing width and scaling depth and fixing depth and scaling width. The precise parameters for each variation are as follows. For our fixed depth experiments, we fix the network depth to 4 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768,1024\}$. The head dimension is fixed to 64. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
  • Hyperparameters: We use a learning rate of $31e-$ 3 with a cosine decay scheduler and weight decay 0.1. We set the maximum training duration to be 10000 steps, with batch size 128 and context length 256.
@@
  • Hyperparameters: We use a learning rate of $41e-$ 4 with a cosine decay scheduler and weight decay 0. We set the maximum training duration to be 30000 steps, with batch size 64 and context length 512.
@@ Appendix B Breakthroughness and Linearity
Figure 8: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 8(a) count and 8(b) addition.
@@ Appendix C Additional Exact Match Accuracy Histogram Plots
Figure 9: Random variation in length generalization (count task). Histograms of exact match accuracy on length $60$ sequences when independently scaling 9(a) width and 9(b) depth.
@@ Appendix D Additional Continuous Metric Plots
We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures. However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt. Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric. We plot the resulting histograms of this metric in Figures 11 and 12 for addition and count respectively. We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
In Figure 13, we plot the analogous KDE plot of the loss-based error metric for the count task. For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition. However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 9 and 12 is still associated with stronger bimodality in the loss plots. For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions. For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
@@
Appendix E CSQA LM Results
E.1 Additional Task
In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA) [43] task and we include the result in Figure 14. Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
@@ E.1 Additional Task
E.2 Impact of Model Size
We repeat the same data mix experiment detailed in Section 2.2 on Qwen2.5-1.5B and report results in Figure 15. In contrast to Figure 7 left, a larger model can learn to perform well on the MMLU task with fewer samples in the training data. For example, Qwen2.5-1.5B can consistently achieve non-trivial performance when the continually pretrained on a data mix that only contains 5% of MMLU auxiliary training samples. In contrast, to achieve a similar performance distribution across 20 runs, Qwen2.5-0.5B requires a data mix that contains 20% of MMLU auxiliary training samples. Interestingly, while the emergence of MMLU capability for Qwen2.5-0.5B showcase bimodal behavior, such a phenomenon is not observed in Qwen2.5-1.5B. As we increase the MMLU ratio in the training data from 0.01% to 0.5%, we instead see a more continuous transition from 0% MMLU performance to non-trivial performance. We qualitatively observe that even with a very small amount of MMLU samples, the last-layer reinitialized Qwen2.5-1.5B model can sometimes learn to output multiple-choice answers but not consistently across all the MMLU test questions. The inconsistent format following behaviors leads to a more continous transition that we observe in Figure 15.
Given that the even the performance distributions with the highest variability in Qwen2.5-1.5B are not clearly bimodal, it appears that bimodality depends on model capacity, at least in our continual pretraining setup. We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model. There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments [3].
Figure 15: Violin plots for last-layer reinitialized Qwen2.5-1.5B models on the MMLU benchmark. The emergence of MMLU capability on Qwen2.5-1.5B shows a more continuous transition.

### v4

@@ 3.2 Emergence Across Data Compositions
We first examine how data composition influences the bimodal capabilities of Qwen2.5-0.5B. Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s performance distribution in ways that mirror the effect of scale.
In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 Raffel et al. (2023), most models (out of 80 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format Hu and Frank (2024). As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($20\%$ MMLU), models across all 80 seeds consistently outperform the 25% random baseline.
@@ 3.3 Emergence Across Model Scales
Like in the synthetic setting, LM model size affects the clustered performance distribution. Using a training data mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 80 seeds and report MMLU test accuracy in Figure 7 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. In fact, larger models reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020). Appendix E.2 shows similar results on the Commonsense QA (Talmor et al., 2019) task.
Figure 8: NLL loss distributions on the MMLU test set. Each distribution is provided as a histogram and as a smoothed KDE. Even with a continuous metric (NLL), we still observe multi-modal distribution across random seeds.
Figure 9: KDE approximation of LM training loss on 10% MMLU mixture. Approximation of the distribution across 80 random seeds is unimodal, smooth, and nearly symmetrical with median close to mean. By contrast, Figure 8 depicts MMLU loss for LMs trained on the same mix as irregularly clustered with multiple modes.
3.4 Is MMLU emergence a mirage?
Is multiple choice performance only bimodal because accuracy is thresholded? As in our synthetic setting (Section 2.4), we must address the effect of the discontinuous metric. In Figure 8, we find that continuous negative log likelihood (NLL) loss metrics remain highly multimodal. The 10% MMLU ratio leads to the most even split between successes and failure in accuracy, and accordingly to clear bimodality in NLL loss, mirroring the link we observed on synthetic tasks. Notably, training data with a 7.5% MMLU ratio produces a performance distribution with a small “success” cluster peaked around NLL=3.4. The 10% MMLU ratio produces a performance distribution with a larger success cluster, also peaking around NLL=3.4. Therefore, the continuous metrics remain roughly bimodal—as is especially clear in the 10% mix setting, which is most evenly split between success and failure.
While recent arguments against emergence (Schaeffer et al., 2024) claim that discontinuities disappear with continuous loss metrics, Figure 8 reveals the opposite effect in random variation—NLL loss reveals clusters that are concealed by accuracy metrics, specifically among low-performance models like those trained on the 5% MMLU mix. (To easily compare these metrics, see Figure 13.)
Is this random clustering particular to the emergent task? As an alternative hypothesis, we might expect all tasks—including training loss—to have similarly multimodal NLL distributions, when approximated with our KDE settings. To the contrary, the pretraining loss produces a smooth, unimodal, and almost symmetric KDE distribution (Figure 9). We therefore conclude that multimodality is linked to emergent tasks—not merely a downstream effect of in-distribution training noise.
4 Discussion
@@
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities. The mode of our performance distributions sharply improves at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 2.5); the transition from the initial unimodal distribution to a bimodal one is sudden.
Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order. Prior studies have documented performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018). Most relevant to us, the high variability of length generalization is well-documented (Zhou et al., 2024b; a). More generally, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution loss. Such differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020). Dodge et al. (2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance. Existing work has also found model runs can cluster in shortcut learning (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation. Our work connects these random clustering effects to the phenomenon of emergence at scale.
Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022). Various proposals have, consequently, aimed to improve length generalization. Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024). Others modify the dataset format by adding scratchpad or Chain-of-Thought (Anil et al., 2022) or by incorporating padding and index hints (Jelassi et al., 2023; Zhou et al., 2024a). Regarding random variation, Zhou et al. (2024a) and Zhou et al. (2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
Emergent abilities of LMs: In LLMs, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020). These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023). Understanding the conditions and mechanisms underlying emergence is a key area of research. Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Srivastava et al., 2023; Schaeffer et al., 2024), some breakthrough capabilities remain emergent. Snell et al. (2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022). For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024). Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. (2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. (2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks. In one surprising result, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how emergent tasks respond to architectural hyperparameter tradeoffs.
@@ Acknowledgments
Reproducibility Statement
Upon publication, we will release the code for both synthetic and LLM experiments, including code, and data. Detailed descriptions of model, data and training are provided in Appendix A, and Section 3.
@@ Appendix A Experimental details on synthetic tasks
Below we provide more details when training decoder-only language models on the count and addition—with index hints—tasks. Hyperparameters are largely adapted from [Zhou et al., 2024a]. We train all of our models to convergence on the train distribution. Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
@@
  • Model scales: As mentioned in Section 2.1, we scale up our models by fixing width and scaling depth and fixing depth and scaling width. The precise parameters for each variation are as follows. For our fixed depth experiments, we fix the network depth to 4 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768,1024\}$. The head dimension is fixed to 64. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
  • Hyperparameters: We use a learning rate of 1e31e-3 with a cosine decay scheduler and weight decay 0.1. We set the maximum training duration to be 10000 steps, with batch size 128 and context length 256.
@@
  • Hyperparameters: We use a learning rate of 1e41e-4 with a cosine decay scheduler and weight decay 0. We set the maximum training duration to be 30000 steps, with batch size 64 and context length 512.
@@ Appendix B Breakthroughness and Linearity
Figure 10: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 10(a) count and 10(b) addition.
@@ Appendix C Additional Exact Match Accuracy Histogram Plots
Figure 11: Random variation in length generalization (count task). Histograms of exact match accuracy on length 6060 sequences when independently scaling 11(a) width and 11(b) depth.
@@ Appendix D Additional Continuous Metric Plots
We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures. However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt. Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric. We plot the resulting histograms of this metric in Figures 14 and 15 for addition and count respectively. We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
In Figure 16, we plot the analogous KDE plot of the loss-based error metric for the count task. For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition. However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 11 and 15 is still associated with stronger bimodality in the loss plots. For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions. For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
Finally, Figure 13 provides a copy of Figure 8 with both accuracy and NLL loss comparable for easy contrast. We note that high-performance clusters are distinguishable in both accuracy and loss, but the thresholded accuracy conceals low-performance clusters which are visible in the continuous distribution.
Figure 13: KDE approximations of performance variation on the MMLU test set. Provided for contrasting accuracy and continuous NLL loss. Top: Approximate distribution of accuracy metrics. Bottom: Approximate distribution of KDE metrics, copied from Figure 8.
@@
Appendix E Additional Language Model Results
E.1 MMLU Results on Qwen2.5-1.5B
We repeat the same MMLU experiment (Section 3.1) on Qwen2.5-1.5B and we include results in Figure 17. Due to computation limit, for each data mixes, we train on 20 random seeds. Similar to Figure 13. models yields the same bimodal distribution in accuracy metrics as well as multi-modal distribution in continuous metric (NLL).
Figure 17: Accuracy and NLL loss for Qwen2.5-1.5B MMLU experiments. We train last-layer reinitialized Qwen2.5-1.5B on data mixes with different ratios of the target task (MMLU). Top: Distribution of accuracy metric. Bottom: Distribution of NLL loss metric.
E.2 CSQA LM Results
In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA, Talmor et al. [2019]) task and we include the result in Figure 18. Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.